OpenStudy (anonymous):

Which of these has the greatest area? Use pi. (3.14) A. rhombus with side lengths: 10 inches, height: 4 inches B. triangle with base length: 12 inches, height: 6 inches C. circle with radius: 4 inches D. trapezoid with base lengths: 3 inches and 11 inches, height: 6 inches

OpenStudy (tkhunny):

Why not calculate the four areas and see?

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Just calculate the areas. Do you have the appropriate formulas?

OpenStudy (anonymous):

No.

OpenStudy (tkhunny):

Perfect. Then you cannot answer this problem. Look them up in a book or online. Write them in here and we can get started.

OpenStudy (anonymous):

1.A=10*4 2.A=(1/2)*12*6 3.A=pi*(4)^2 4.A={(3+11)/2}*6 calculate